# Unpacking Positional Encoding in Transformers: A Spectral Analysis of Content-Position Coupling

### v1

@@ 6 Conclusion and Limitations
In conclusion, we have introduced a spectral-theoretic framework that connects positional encoding designs to the eigenstructure of attention logits and shown—through synthetic tasks and head-wise ablations—that these spectral signatures reliably predict learning dynamics and internal processing patterns (e.g., RoPE’s “single-head deposit”). Our results illuminate how different coupling mechanisms shape the interplay between content and position in attention and suggest practical strategies—such as controlled mixing—to achieve both spectral stability and distributed encoding. These insights lay the groundwork for principled design of more robust and interpretable PE schemes in Transformer models.
Limitations: This study is primarily theoretical and relies on controlled synthetic tasks, so its findings may not directly translate to performance on complex, real-world datasets or broad benchmarks. We also explore only a limited range of task complexities and positional relationships; more intricate challenges—such as sequence reversal or Dyck-language recognition—likely exceed a Transformer’s positional reasoning capabilities. Finally, our proposed connection between single-head deposit patterns and length extrapolation remains theoretical and requires direct empirical validation.
@@ Appendix A Notations List
Appendix B Detailed Theoretical Analysis
@@
Our analysis is predicated on a core theoretical link regarding the role of spectral properties in learning. We rely on the established principle that favorable spectral properties of matrices involved in optimization—such as a more contracted eigenvalue spectrum (tighter bounds) or a smaller spectral norm—correlate with faster and more stable convergence of learning algorithms. Building upon this, our theoretical framework posits that the rapid and stable convergence of the positional coupling function, driven by these favorable spectral properties, will lead to a specific phenomenon during training: the localization of this function within a limited set of attention heads in shallow layers, resulting in the empirically observed deposit pattern [30].
@@
We use the fomula in [3] that a logit value of i and j in $\mathbf{L}_{\text{RoPE\_complex}}$ is
@@ 2. RoPE
It is easy to verify that the solution space of RoPE is much larger than that of Relative Position Encoding. This can be considered from the following two aspects. First, the expression of RoPE can make the key-value pairs with a fixed distance meet the requirements as long as the angle distribution meets the requirements, while the expression of Relative Position Encoding requires a more stringent angle distribution because the bias has no significant effect. Second, RoPE actually only needs a significant angle to easily meet the conditions, which actually corresponds to the cause of Massive Value described in  [22].
@@
However, a strict comparison requires citation of Szegő’s Theorem on Eigenvalue Distribution [25]. So let us calculate the generating function (sign function) corresponding to $W\circ E$ and $W$ respectively.
@@
[…]
@@
[…]
@@
The first equality forces **all terms in the sum for $P$ to share a common phase**, i.e. $e^{\mathrm{i}(\theta_{i}+i\theta^{\star})}$ is the same for every $i$; the second forces this common phase to be 00 or $\pi$, which does not affect alignment. Thus
@@
Appendix C Support Experiment
C.1 Experiment Setup
@@
C.2 Experiments of Head Ablation on Other PE Methods
@@
The similarities among NoPE, Relative PE, and ALiBi are consistent with previous studies on how NoPE obtains position information through causal masks [5, 20].
C.3 Some Experiments on Partial RoPE
@@
C.4 Why Position Embedding Complex?
@@
Appendix D Relationship to Length Generation
@@
More than one study has shown that transformer’s understanding of position information starts from the first layer [19, 9, 5]. The key to successful length generalization is that responses to relative position information within the training length can be naturally transferred to a longer length (test length). Therefore, the lower the coupling between location and content, the better the scalability. If the coupling degree is high, it is necessary to learn the coupling mode.
Our experimental results show that the deposit pattern is not difficult to improve through dropout, which means that once a head of trasnformer of RoPE captures the pattern of training position information, the other heads no longer capture the position information. Therefore, the number of parameters responding to position content coupling is much less than that of relative position encoding and no position encoding, resulting in higher compression and lower regularization.
@@
At the same time, Deposit Patterns shows that only some frequency bands are fully trained during training like [4]. These frequency bands are passed through the fully connected layer, spreading the influence to most of the frequency bands in the deep layer, which means that in fact only specific relative position differences are considered by the training process, so the length generalization ability is drastically affected.

### v2

@@ 7 Conclusion and Limitations
In this work, we introduced a novel framework that leverages spectral theory to deconstruct how positional encoding (PE) mechanisms operate within the Transformer attention matrix. We demonstrated that this framework provides a powerful lens, allowing us to predict learning dynamics and uncover previously hidden processing patterns. Our primary discovery, the "single-head deposit pattern" in RoPE, serves as a clear mechanistic explanation for the observed paradoxes in its performance. By identifying the distinct signatures of additive and multiplicative coupling, our analysis not only diagnoses the cause of this over-specialization but also points toward a solution. The success of a hybrid architecture validates our theory and suggests that controlled mixing of positional signals is a promising strategy for achieving both spectral stability and robust, distributed representations. Furthermore, we exploit gradient flow to verify that this is an intrinsic property of the RoPE.
Limitations: Firstly, the connection we hypothesize between the deposit pattern and RoPE’s known struggles with length extrapolation requires direct empirical validation. Secondly, future work should probe the limits of these positional mechanisms on more complex, algorithmic tasks such as sequence reversal or Dyck language recognition. However, most of these tasks are beyond the capabilities of transformers, so more complex controlled analysis is a serious challenge for future work.
Acknowledgments
This work was supported by the National Natural Science Foundation of China under Grant 62372448.
Acknowledgment of LLM Usage
During the preparation of this manuscript, large language models (LLMs) were employed in a limited and auxiliary capacity. Specifically, their usage was restricted to the following three aspects: (1) checking grammar and expression at the sentence level, thereby providing local linguistic refinement; (2) performing global polishing after the draft was completed, ensuring that the overall exposition conforms to idiomatic English usage; and (3) improving the readability of the proof details presented in the appendix.
At no stage were LLMs used for generating research ideas, developing arguments, or modifying the substantive content of this work. Their sole role was to assist in enhancing clarity and effectiveness of communication.
@@ Appendix A Notations List
Appendix B Detailed Method Analysis
@@
Our analysis is predicated on a core theoretical link regarding the role of spectral properties in learning. We rely on the established principle that favorable spectral properties of matrices involved in optimization—such as a more contracted eigenvalue spectrum (tighter bounds) or a smaller spectral norm—correlate with faster and more stable convergence of learning algorithms. Building upon this, our theoretical framework posits that the rapid and stable convergence of the positional coupling function, driven by these favorable spectral properties, will lead to a specific phenomenon during training: the localization of this function within a limited set of attention heads in shallow layers, resulting in the empirically observed deposit pattern Boyd and Vandenberghe (2004).
@@
We use the fomula in Su et al. (2024) that a logit value of i and j in $\mathbf{L}_{\text{RoPE\_complex}}$ is
@@ 2. RoPE
It is easy to verify that the solution space of RoPE is much larger than that of Relative Position Encoding. This can be considered from the following two aspects. First, the expression of RoPE can make the key-value pairs with a fixed distance meet the requirements as long as the angle distribution meets the requirements, while the expression of Relative Position Encoding requires a more stringent angle distribution because the bias has no significant effect. Second, RoPE actually only needs a significant angle to easily meet the conditions, which actually corresponds to the cause of Massive Value described in  Jin et al. (2025).
The emergence of the deposit pattern may share a common theoretical origin: the spectral properties of the attention logit matrix. Gradient-based optimization is known to be more stable for matrices with tightly clustered eigenvalues (Boyd and Vandenberghe, 2004). As suggested by Szegő’s theorem (Grenander, 1958), the multiplicative interaction in RoPE is theorized to contract the eigenvalue spectrum of the positional component more effectively than additive methods.
@@
The structure of the attention logit matrix $\mathbf{L}$, particularly the properties of its Toeplitz or Toeplitz-like components, provides crucial insights into the stability and dynamics of positional information processing. The width of the eigenvalue spectrum often has a great impact on the optimization process.
Theorem B.1 (Szegő’s Theorem on Eigenvalue Distribution (From Grenander (1958))).
@@
However, a strict comparison requires citation of Szegő’s Theorem B.1 on Eigenvalue Distribution Grenander (1958). So let us calculate the generating function (sign function) corresponding to $W\circ E$ and $W$ respectively.
@@
$\max_{\theta\in\mathbb{R}}\!\bigl|f(\theta)\bigr|\;\leq\;\max_{\theta\in\mathbb{R}}\!\bigl|g(\theta)\bigr|\;=\;2\sum_{i=0}^{N-1}w_{i}.$
@@
$|f(\theta)|\;=\;2\bigl|\Re P(e^{\mathrm{i}\theta})\bigr|\;\leq\;2\!\left|P(e^{\mathrm{i}\theta})\right|\;\leq\;2\sum_{i=0}^{N-1}w_{i}\;=\;\bigl|g(0)\bigr|.$
@@
The first equality forces **all terms in the sum for $P$ to share a common phase**, i.e. $e^{\mathrm{i}(\theta_{i}+i\theta^{\star})}$ is the same for every $i$; the second forces this common phase to be $0$ or $\pi$, which does not affect alignment. Thus
@@
Appendix C Detailed Proofs for Theoretical Analysis
This appendix provides the detailed proofs for the propositions and theorems presented in Section 6.
C.1 Preliminaries and Notation
We consider a Transformer with $L$ layers and $H$ heads. For a batch of $N$ samples, $(i_{s},j_{s})$ are the token positions for sample $s$, and $d_{s}=|j_{s}-i_{s}|$ is the relative distance. The empirical measure of distances is $\hat{\mu}_{N}(d):=\frac{1}{N}\sum_{s=1}^{N}\mathbf{1}\{|j_{s}-i_{s}|=d\}$. The attention mechanism is defined by $S^{(h,l)}=Q^{(h,l)}K^{(h,l)\top}/\sqrt{d_{h}}$, $A^{(h,l)}=\text{softmax}_{\text{row}}(S^{(h,l)})$, and $Y^{(h,l)}=A^{(h,l)}V^{(h,l)}$. We use the following constant, derived from bounded operator norms and loss gradients:
$C^{(h)}:=\|W_{O}\|_{2}\,\|V^{(h,L)}\|_{2}\,\sup_{s}\|\partial\ell_{s}/\partial y\|_{2}.$ (6)
We also denote the row sharpness as $\chi_{L}:=\max_{i}\|A^{(h,L)}_{i,:}\|_{2}$ and the minimum anchor attention as $a_{*}^{(h)}(d):=\inf_{s:\,d_{s}=d}A^{(h,L)}_{u_{s},j_{s}}$.
C.2 Softmax Gradient Calculus
Lemma C.1 (Row-softmax Gradient Identity).
For any layer $l$, head $h$, sample $s$, and row $i$,
$\frac{\partial\ell_{s}}{\partial S^{(h,l)}_{i,j}}=A^{(h,l)}_{i,j}\Big(\Lambda^{(h,l)}_{i,j}-\langle\Lambda^{(h,l)}_{i,:},A^{(h,l)}_{i,:}\rangle\Big)$, (7)
where $\Lambda^{(h,l)}:=(\partial\ell_{s}/\partial Y^{(h,l)})V^{(h,l)\top}$.
Proof.
The result follows from the chain rule. The Jacobian of a row-wise softmax is $J_{i}=\mathrm{diag}(A_{i,:})-A_{i,:}A_{i,:}^{\top}$. The upstream gradient is $\partial\ell/\partial A=(\partial\ell/\partial Y)V^{\top}$. Combining these yields the identity. ∎
Lemma C.2 (Quantitative Anchor Gain).
Let $j^{\star}$ be an anchor column for row $i$. Let $J_{i}$ be the softmax Jacobian for row $i$. For the canonical basis vector $e_{j^{\star}}$, we have:
$\|P_{e_{j^{\star}}}\,J_{i}^{\top}e_{j^{\star}}\|_{2}=A_{i,j^{\star}}(1-A_{i,j^{\star}})\quad\text{and}\quad\|P_{(e_{j^{\star}})^{\perp}}\,J_{i}^{\top}e_{j^{\star}}\|_{2}\leq A_{i,j^{\star}}(1-A_{i,j^{\star}}).$ (8)
Proof.
A direct computation gives $J_{i}^{\top}e_{j^{\star}}=A_{i,j^{\star}}e_{j^{\star}}-A_{i,j^{\star}}A_{i,:}$. The component along $e_{j^{\star}}$ is $A_{i,j^{\star}}(1-A_{i,j^{\star}})$. The squared norm of the orthogonal component is $A_{i,j^{\star}}^{2}\sum_{k\neq j^{\star}}A_{i,k}^{2}\leq A_{i,j^{\star}}^{2}(\sum_{k\neq j^{\star}}A_{i,k})^{2}=A_{i,j^{\star}}^{2}(1-A_{i,j^{\star}})^{2}$. Taking square roots yields the result. ∎
C.3 Proof of Proposition 6.1 (RoPE Top-Layer Seed)
Proof.
From Lemma C.1, for a sample $s$ with anchor $u_{s}$ and target $j_{s}$:
$\frac{\partial\ell_{s}}{\partial S^{(h,L)}_{u_{s},j_{s}}}=A^{(h,L)}_{u_{s},j_{s}}\Big(\Lambda^{(h,L)}_{u_{s},j_{s}}-\langle\Lambda^{(h,L)}_{u_{s},:},A^{(h,L)}_{u_{s},:}\rangle\Big)$. (9)
By Assumption (A3), the first term $\Lambda^{(h,L)}_{u_{s},j_{s}}=\langle\partial\ell_{s}/\partial Y^{(h,L)}_{u_{s},:},V^{(h,L)}_{j_{s},:}\rangle\geq\eta_{s}^{(h)}$. For the second term, Cauchy-Schwarz and the definition of $C^{(h)}$ yield:
$|\langle\Lambda_{u_{s},:},A_{u_{s},:}\rangle|\leq\|\Lambda_{u_{s},:}\|_{2}\,\|A_{u_{s},:}\|_{2}\leq\|\partial\ell_{s}/\partial Y_{u_{s},:}\|_{2}\,\|V^{(h,L)}\|_{2}\,\chi_{L}\leq C^{(h)}\chi_{L}.$ (10)
Combining these, and using $A^{(h,L)}_{u_{s},j_{s}}\leq 1$:
$\frac{\partial\ell_{s}}{\partial S^{(h,L)}_{u_{s},j_{s}}}\geq A^{(h,L)}_{u_{s},j_{s}}\eta_{s}^{(h)}-A^{(h,L)}_{u_{s},j_{s}}C^{(h)}\chi_{L}\geq a_{*}^{(h)}(d_{s})\eta_{*}^{(h)}(d_{s})-C^{(h)}\chi_{L}.$ (11)
Summing over all samples $s$ with $d_{s}=d$ to get $H^{(h)}_{L}(d)$ gives the result:
$H^{(h)}_{L}(d)\geq N\,\hat{\mu}_{N}(d)\,\Big(a_{*}^{(h)}(d)\,\eta_{*}^{(h)}(d)-C^{(h)}\chi_{L}\Big).$ (12)
This is strictly positive if the learning signal term $a_{*}\eta_{*}$ outweighs the interference term $C\chi_{L}$. ∎
C.4 Proof of Proposition 6.2 (ALiBi Cancellation)
Proof.
We provide a constructive proof. Fix a distance bucket $d$ and partition its samples $\mathcal{S}_{d}$ into pairs $(s,s^{\prime})$. For each pair, we construct their token embeddings to achieve cancellation.
  1. 1.
    Identical Attention Matrix: By (A3), choose $t$ such that $W_{Q}t=W_{K}t=0$ and $W_{V}t\neq 0$. For a base token embedding $x_{0}$, set the anchor token embedding for sample $s$ to be $x_{s}=x_{0}+t$ and for $s^{\prime}$ to be $x_{s^{\prime}}=x_{0}-t$. All other token embeddings are identical for $s$ and $s^{\prime}$. Since the query and key projections of $t$ are zero, all dot products $q_{i}\cdot k_{j}$ are identical for both samples. As ALiBi adds the same bias $b_{h}(j-i)$, the scores $S^{(h,L)}$ and attention matrices $A^{(h,L)}$ are identical for $s$ and $s^{\prime}$.
  2. 2.
    Flipped Gradient Signal: Since $W_{V}t\neq 0$, the value vector at the anchor row $u_{s}$ is flipped: $V^{(h,L)}_{u_{s},:}(s^{\prime})\approx V^{(h,L)}_{u_{s},:}(s)-2W_{V}t$. By appropriate choice of $x_{0}$, this can be made an exact sign flip. This implies that the term $\Lambda^{(h,L)}_{u_{s},:}(s^{\prime})=-\Lambda^{(h,L)}_{u_{s},:}(s)$.
  3. 3.
    Pairwise Cancellation: Applying Lemma C.1 to the anchor edge $(u_{s},j_{s})$ for both samples, we find their respective gradients are equal and opposite, summing to zero. Repeating this for all pairs makes $H^{(h)}_{L}(d)=0$.
C.5 Proof of Theorem 6.1 (Exponential Amplification)
Let $\mathcal{U}$ be the subspace spanned by gradient directions corresponding to the seed-positive buckets. Let $g^{(l)}$ be the gradient vector at layer $l$, so $g^{(l)}=J^{(l)\top}g^{(l+1)}$.
Lemma C.3 (Layerwise Directional Advantage).
Under the assumptions, there exists $\beta_{l}\geq 0$ such that for any $v\neq 0$,
$\frac{\|P_{\mathcal{U}}J^{(l)\top}v\|_{2}}{\|P_{\mathcal{U}^{\perp}}J^{(l)\top}v\|_{2}}\geq\frac{1-\alpha_{l}}{1+\alpha_{l}}(1+\beta_{l}).$ (13)
The gain $\beta_{l}$ is lower-bounded by a term proportional to the summed strength of the positive seeds from Proposition 6.1.
Proof.
The Jacobian $J^{(l)}$ is decomposed into the within-attention part $\mathcal{O}^{(l)}$ and the rest of the block $\mathcal{R}^{(l)}$. By (A1), the $\mathcal{R}^{(l)}$ part contributes the factor $(1-\alpha_{l})/(1+\alpha_{l})$. The gain $\beta_{l}$ arises from $\mathcal{O}^{(l)}$. Lemma C.2 establishes a directional gain for the anchor component within the softmax Jacobian. When composed with the value projection and the upstream gradient from (A2), the anchor direction receives a coherent signal. Aggregating over the anchor rows in subspace $\mathcal{U}$ yields a net directional advantage, which extends from the basis vectors of $\mathcal{U}$ to any vector $v$ by a convexity argument. ∎
Proof of Theorem 6.1.
Applying Lemma C.3 to the backpropagation recurrence $g^{(l)}=J^{(l)\top}g^{(l+1)}$:
$\mathrm{SNR}_{l}=\frac{\|P_{\mathcal{U}}J^{(l)\top}g^{(l+1)}\|_{2}}{\|P_{\mathcal{U}^{\perp}}J^{(l)\top}g^{(l+1)}\|_{2}}\geq\gamma_{l}\cdot\frac{\|P_{\mathcal{U}}g^{(l+1)}\|_{2}}{\|P_{\mathcal{U}^{\perp}}g^{(l+1)}\|_{2}}=\gamma_{l}\cdot\mathrm{SNR}_{l+1}.$ (14)
Iterating this inequality from layer $L-1$ down to $l$ yields the exponential product. The same multiplicative logic applies to the vector components corresponding to the top two heads within the subspace $\mathcal{U}$, proving margin amplification. ∎
C.6 Remark: Conceptual Link to Convolution Kernels
It is helpful to conceptually frame the aggregated gradient $H^{(h)}_{L}(d)$ as a discrete convolution. If we define the empirical distribution of distances as a signal $\hat{\mu}_{N}$, and the expected gradient contribution for a given relative distance $\Delta$ as a "deposit kernel" $\kappa^{(h)}(\Delta)$, then the total aggregated gradient is their convolution: $H^{(h)}_{L}=N(\hat{\mu}_{N}*\kappa^{(h)})$.
From this perspective, our proof demonstrates a key difference in the structure of these implicit kernels:
  • For RoPE, the kernel $\kappa^{(h)}(\Delta)$ has a trigonometric structure due to the rotational mechanism. This structure is analogous to a positive-definite kernel, which resists being driven to zero and ensures a positive "seed" is deposited.
  • For ALiBi, the kernel $\kappa^{(h)}(\Delta)$ is merely affine (linear plus a constant). This simple structure allows for exact cancellation when convolved with a symmetric distance distribution, explaining why no seed is guaranteed.
This intuitive framing aligns with the rigorous proof and reinforces the conclusion that the deposit pattern is an inherent property of RoPE’s multiplicative design.
Appendix D Support Experiment
D.1 Experiment Setup
@@
D.2 Complete Visualization of Layer-by-Layer Head Ablation
For completeness, we provide the full layer-by-layer head ablation results for the RoPE models trained on Task 1. These figures expand upon the summary violin plots in Fig. 4, showing the per-head accuracy after zeroing out each attention head at every layer.
How to read the plots.
Each subplot corresponds to one Transformer layer, and each point plots the test accuracy after ablating a single head. Large vertical drops indicate heads whose removal severely harms performance, revealing where positional reasoning is concentrated. Flat curves indicate heads that contribute minimally and whose ablation produces no measurable degradation.
6-layer RoPE model.
Figure 9 shows the full ablation for the 6-layer model. The first layer contains a single dominant head whose ablation causes a large accuracy collapse, while all remaining heads exhibit near-flat profiles across layers. This fine-grained visualization matches the condensed “deposit pattern” presented in the main text.
Figure 9: Full head-ablation curves for the 6-layer RoPE model on Task 1. Each subplot shows the test accuracy after ablating a single attention head in the corresponding layer. A single head in Layer 1 produces a catastrophic accuracy drop, while all other heads across all layers induce only minor changes. This confirms that RoPE concentrates positional reasoning into one early-layer head, even in shallower architectures.
8-layer RoPE model.
Figure 10 reports the same analysis for an 8-layer RoPE model. Despite the increased depth, the pattern is qualitatively identical: a single early-layer head carries almost all positional reasoning, and deeper layers show only small, noise-level fluctuations across heads. This demonstrates that the deposit pattern is not sensitive to depth and remains stable across architectures.
Figure 10: Full head-ablation curves for the 8-layer RoPE model on Task 1. The same localization pattern reappears at larger depth: a unique critical head in Layer 1 causes a large accuracy collapse, and all remaining heads across Layers 2–8 have negligible impact. The persistence of this topology across depth demonstrates the stability and robustness of the RoPE deposit pattern.
Summary.
These complete ablation maps confirm that the deposit pattern is a robust and highly localized effect: RoPE consistently routes positional reasoning into one head in the earliest layers, with all other heads acting effectively position-agnostic for this task.
D.3 Experiments of Head Ablation on Other PE Methods
@@
The similarities among NoPE, Relative PE, and ALiBi are consistent with previous studies on how NoPE obtains position information through causal masks Kazemnejad et al. (2023); Haviv et al. (2022).
D.4 Some Experiments on Partial RoPE
@@
D.5 Why Position Embedding Complex?
@@
Appendix E Relationship to Length Generation
@@
Why the deposit pattern reflects a training bias rather than modularity.
Although the deposit pattern superficially resembles a form of head-level modularity, our evidence suggests that it is instead a training-induced bias specific to RoPE’s multiplicative structure. First, multiplicative PE suppresses activation-level positional directions (Section 5.4), causing early-layer heads to compete for the positional signal. Once a single head captures the content–position interaction, RoPE’s strong inductive bias quickly eliminates positional gradients for the other heads, preventing them from acquiring similar capabilities. This is fundamentally different from deliberate modularization: the specialization emerges not because the model decomposes the task across heads, but because the training dynamics collapse positional reasoning into the earliest head that happens to align with the RoPE kernel.
Second, the deposit head does not learn a general positional module; it learns a narrow set of frequency bands that are amplified by RoPE’s rotation mechanism. As noted in Ermo,Hua et al. (2025), only a subset of frequencies is sufficiently trained, and their influence is later propagated through feed-forward mixing. This leads to a compressed, low-rank representation of positional structure, which is efficient for the training distribution but fragile outside it.
Taken together, the deposit pattern is better interpreted as a training bias toward early collapse of positional responsibility, rather than evidence for a stable or interpretable modular decomposition of positional reasoning. RoPE’s multiplicative interaction creates a “winner-takes-all” dynamic, where one head monopolizes positional information not because the architecture encourages modularity, but because the optimization process and frequency structure favor such collapse.
Further evidence from tasks beyond pairwise position reasoning.
RoPE enables efficient pairwise position–content coupling but does not equip the model with the compositional machinery needed for length-generalizing tasks that require combining several positional relations. To test whether the deposit head learned by RoPE represents a reusable “positional module,” we designed a family of tasks requiring the model to combine multiple pairwise distances (e.g., predicting the sum of several trigger-word distances). Despite the task being a straightforward extension of our synthetic setup, all RoPE-based models failed to generalize beyond the training length, even when the training accuracy was near-perfect. In contrast to the pairwise distance task—where positional responsibility collapses into a single head—the multi-trigger task requires the model to compose several independent positional relations, a capability that is architecturally challenging for standard Transformers.
This observation aligns with prior findings on tasks such as Dyck languages, context-free composition, and stack-like dependencies (e.g., Chi et al. (2023)): Transformers struggle when a task requires retrieving, combining, or manipulating multiple positional relations simultaneously. The failure of RoPE to generalize in our multi-trigger setting therefore does not indicate a lack of training or capacity, but rather reveals that the single-head deposit behavior does not constitute a modular or compositional positional representation. Instead, the deposit head encodes a narrow, task-specific mapping tied to the training distribution, and—unlike a true module—it cannot be recombined or composed when the task demands multiple positional deductions.
Taken together, these results reinforce that the deposit pattern reflects a training-induced collapse of positional responsibility rather than the emergence of a reusable module. RoPE enables efficient pairwise position–content coupling but does not equip the model with the compositional machinery needed for length-generalizing tasks that require combining several positional relations.
@@
More than one study has shown that transformer’s understanding of position information starts from the first layer Zuo et al. (2025); Chi et al. (2023); Kazemnejad et al. (2023). The key to successful length generalization is that responses to relative position information within the training length can be naturally transferred to a longer length (test length). Therefore, the lower the coupling between location and content, the better the scalability. If the coupling degree is high, it is necessary to learn the coupling mode. At the same time, Deposit Patterns shows that only some frequency bands are fully trained during training like Ermo,Hua et al. (2025). These frequency bands are passed through the fully connected layer, spreading the influence to most of the frequency bands in the deep layer, which means that in fact only specific relative position differences are considered by the training process, so the length generalization ability is drastically affected.
Relationship with massive value in RoPE.
Recent empirical studies have reported that Transformers trained with RoPE often develop a small number of extremely high-norm rows in $W_{Q}$ and $W_{K}$ (Jin et al., 2025), a behavior not observed under additive positional encodings. Although this phenomenon is not the focus of our work, the Toeplitz formulation naturally explains its underlying mechanism.
RoPE mixes content and position multiplicatively through a frequency-indexed Toeplitz kernel $G_{\mathbf{e}}$. In our formulation, each logit entry takes the form
$\langle q_{i},k_{j}\rangle_{\mathrm{RoPE}}=\langle q_{i},k_{j}\rangle\circ(G_{\mathbf{e}})_{i-j},$
where $(G_{\mathbf{e}})_{i-j}$ is a complex rotation encoding relative displacement. Importantly, each frequency band of RoPE corresponds to a fixed Toeplitz diagonal, and gradients flowing through that diagonal are accumulated entirely on the rows of $W_{Q},W_{K}$ associated with the same frequency.
This creates a “frequency bottleneck”:
  • if a particular diagonal of the Toeplitz kernel becomes predictive early in training,
  • then its associated frequency bands receive disproportionately large gradient flow,
  • causing the corresponding rows of $W_{Q},W_{K}$ to grow rapidly in norm,
  • while other frequencies remain under-trained.
Thus the massive-value rows arise as a direct consequence of gradient concentration onto a small subset of Toeplitz diagonals, not from instability of the optimization process. Additive PEs do not exhibit this effect because their position signal enters only through additive $p_{i}$–based cross-terms, so no frequency-indexed diagonals receive concentrated multiplicative amplification.
Under our unified framework, the massive-value phenomenon is therefore a natural by-product of RoPE’s multiplicative structure, in which positional information is stored directly in selected frequency bands of $W_{Q},W_{K}$ rather than in explicit activation-level positional vectors.